The joint density function of the random variables X, Y is given as

f(x, y) = { 0.4 (2x + 3y), 0 < x < 1, 0 < y< 1

0, elsewhere

1)Determine P [ 0.33 < X < 2/3 |Y > 3 /4 ]


1
Expert's answer
2022-05-12T09:22:22-0400

1). Check, whether "f(x,y)" is a valid joint density function: "\int_0^1\int_0^10.4(2x+3y)dydx=0.4\int_0^1(2xy+\frac32y^2)|_0^1dx=0.4\int_0^1(2x+\frac32)dx=0.4(x^2+\frac32x)|_0^1=1"Thus, it is a valid joint density function.

Remind the formula: "P(A|B)=\frac{P(A\cap B)}{P(B)}". From the latter we get: "P(0.33<X<\frac23|Y>\frac34)=\frac{P(0.33<X<\frac23,Y>\frac34)}{P(Y>\frac34)}".

Consider separately each probability in the last formula: "P(0.33<X<\frac23,Y>\frac34)=0.4\int_{0.33}^{\frac23}\int_{\frac34}^{1}(2x+3y)dydx=0.4\int_{0.33}^{\frac23}(2xy+\frac32y^2)_{\frac34}^{1}dx=0.4\int_{0.33}^{\frac23}(2x+\frac32-\frac32x-\frac{27}{32})dx=0.4\int_{0.33}^{\frac23}(\frac12x+\frac{21}{32})dx=0.4(\frac14x^2+\frac{21}{32}x)_{0.33}^{\frac23}\approx0.1219"

The latter value is rounded to "4" decimal places.

"P(Y>\frac34)=0.4\int_0^1\int_{\frac34}^1(2x+3y)dydx=0.4\int_0^1(2xy+\frac32y^2)|_{\frac34}^1dx=0.4\int_0^1(2x+\frac32-\frac32x-\frac{27}{32})dx=0.4\int_0^1(\frac12x+\frac{21}{32})dx=0.4\left(\frac{1}{4}x^2+\frac{21}{32}x\right)|_0^1=0.4\left(\frac14+\frac{21}{32}\right)=\frac{29}{80}"

Thus, the probability is the following: "P(0.33<X<\frac23|Y>\frac34)\approx0.1219\cdot\frac{80}{29}\approx0.336"

Answer: "P(\frac13<X<\frac23,Y>\frac34)\approx0.336" (it is rounded to "3" decimal places)